What volume in ml of 0.8 M HCl is needed to completely react with 91 grams of the CaCO3?


1
Expert's answer
2021-04-06T04:33:58-0400

2HCl+CaCO3=CaCl2+H2O+CO2


n (CaCO3)="m(CaCO3)\/M(CaCO3)" "=" "91 g\/100 g\/mol" = 0.91 mol

n(HCl)="2*n(CaCO3)" "=" "2*0.91 mol= 1.82 mol"

V(HCl)="n(HCl)\/c(HCl)" = "1.82 mol\/ 0.8 mol\/L" = 2.275 L= 2275 ml


Answer: 2275 ml